OpenStudy (anonymous):

How do I find the domain of f(x, y) =4-x^2-2y^2

OpenStudy (anonymous):

\[\large f(x, y) =4-x^{2}-2y^{2} \]

OpenStudy (anonymous):

i know how to do it when there are fractions and the denominator cannot equal 0 or the natural logarithm of something most be greater than zero but dont have any restrictions on this one so not sure what to do. @hartnn

OpenStudy (anonymous):

Perhaps the simplest way is to consider a 2D section (contour map). So u set f(x,y) = constant c (if c is 0 y have an ellipse, then vary c up and down, see what happens)

hartnn (hartnn):

i would say since x is not restriced because of any square root sign or denominator, the domain is all real numbers.....

hartnn (hartnn):

*restricted

OpenStudy (zarkon):

this is a polynomial. the domain for a polynomial is not restricted. Therefore the domain is \(\mathbb{R}\times\mathbb{R}\)